# Unsupervised Contrastive Analysis for Salient Pattern Detection using Conditional Diffusion Models

### v1

@@ Abstract
Contrastive Analysis (CA) regards the problem of identifying patterns in images that allow distinguishing between a background (BG) dataset (i.e. healthy subjects) and a target (TG) dataset (i.e. unhealthy subjects). Recent works on this topic rely on variational autoencoders (VAE) or contrastive learning strategies to learn the patterns that separate TG samples from BG samples in a supervised manner. However, the dependency on target (unhealthy) samples can be challenging in medical scenarios due to their limited availability. Also, the blurred reconstructions of VAEs lack utility and interpretability. In this work, we redefine the CA task by employing a self-supervised contrastive encoder to learn a latent representation encoding only common patterns from input images, using samples exclusively from the BG dataset during training, and approximating the distribution of the target patterns by leveraging data augmentation techniques. Subsequently, we exploit state-of-the-art generative methods, i.e. diffusion models, conditioned on the learned latent representation to produce a realistic (healthy) version of the input image encoding solely the common patterns. Thorough validation on a facial image dataset and experiments across three brain MRI datasets demonstrate that conditioning the generative process of state-of-the-art generative methods with the latent representation from our self-supervised contrastive encoder yields improvements in the generated image quality and in the accuracy of image classification. The code is available at https://github.com/CristianoPatricio/unsupervised-contrastive-cond-diff.
@@ 1 Introduction
Figure 1: Overview of the proposed contrastive-guided conditional diffusion model. The common features encoder $f_{\theta}(.)$ receives the input image $x_{0}$ to encode the latent representation $z_{bg}$, which captures common information. The decoder $p_{\theta}(x_{0}|x_{t},z_{bg})$ receives the noisy sample $x_{t}$ jointly with the common feature $z_{bg}$ to reconstruct the normal version of the input image. The saliency map is obtained by $\Delta_{SM}=\left|x_{0}-\hat{x_{0}}\right|$, where high values indicate the salient patterns. Intuitively, the class of the image can be determined based on the magnitude of $\bar{\Delta_{SM}}$.
@@
Despite substantial progress in image classification models, they still face two significant challenges: i) dependence on extensive amounts of labelled data; and ii) limited interpretability. These challenges are particularly critical in sectors such as healthcare, where explaining the decision process is essential for clinicians to trust the model outcome, and unhealthy samples are scarce and hard to obtain Patrício et al. (2023). Contrastive Analysis (CA) represents a promising solution to this problem Zou et al. (2013); Abid et al. (2018); Weinberger et al. (2022) by learning the fundamental generative factors that distinguish a target (TG) dataset from a background (BG) dataset (referred to as salient patterns) and are shared across both datasets (referred to as common patterns). However, state-of-the-art CA methodologies have several problems: i) they often result in blurred images when utilizing Variational Autoencoders (VAEs) or are susceptible to mode collapse and unstable training when using Generative Adversarial Networks (GANs) Schlegl et al. (2019); Carton et al. (2024); ii) they depend on the availability of both healthy and unhealthy samples during training; iii) they struggle to preserve common patterns of the original image during the reconstruction process.
To address these problems, this work redefines the CA task by employing a self-supervised contrastive encoder (referred to as common features encoder) to learn a latent representation, using samples exclusively from the BG dataset during training, and approximating the salient variation factor by leveraging data augmentation techniques such as random cutout or Gaussian noise (e.g. presence of eyeglasses in CelebA or presence of tumor in brain MRIs). The problems of previous generative methods are addressed by leveraging Denoising Diffusion Probabilistic Models (DDPM) Ho et al. (2020) to learn the distribution of the BG dataset (control group) conditioned on the learned latent representation by the common features encoder. During inference, the diffusion model processes unseen samples from BG or TG distributions by substituting the salient patterns with common patterns learned during training. This enables the classification of the input image as background (normal) or target (anomalous) by assessing the magnitude of the reconstruction error.
Extensive validation on facial imagery data demonstrates improved similarity between original and reconstructed images. Additionally, comparative analysis with state-of-the-art methods for salient pattern detection for brain MRI further supports the efficacy of our approach. Our major contributions are as follows: i) a methodology for learning shared common information between background and target distributions, thus allowing the generation of a healthy (normal) version of the input image by encoding only its common information; ii) a common features encoder capable of learning input representations that are both class-invariant and instance-aligned. iii) a comprehensive evaluation of our approach on a facial imagery dataset and three brain MRI datasets, encompassing both healthy (background) and tumor (target) images.
@@ 2 Related Work
Early work in contrastive analysis (CA) relied on the use of Variational Autoencoders (VAEs) Baur et al. (2021); Behrendt et al. (2022); Louiset et al. (2023). A recent method proposed by Louiset et al. (2023), SepVAE, aims to distinguish common (healthy) from class-specific (unhealthy) patterns in image data. They utilized VAEs with regularization to encourage disentanglement between common and salient representations, along with a classification loss to separate target and background salient factors. However, the resulting reconstructions were often blurry, limiting their interpretability and utility. Alternative approaches Schlegl et al. (2019); Carton et al. (2024) leveraging Generative Adversarial Networks (GANs) have been explored, but they suffer from issues like mode collapse and unstable training. More recently, Diffusion Denoising Probabilistic Models Song et al. (2021); Preechakul et al. (2022) (DDPMs) have emerged as a promising alternative for high-quality image generation, addressing the drawbacks of both GANs and VAEs. Our work aligns closely with the setting of salient pattern detection. In the domain of brain MRI, Behrendt et al. (2024) introduced a patch-based diffusion model for discovering salient patterns in an unsupervised manner in brain MRIs. They divide the input image into predefined patches and apply noise to each patch individually in the forward pass. In the backward pass, the partly noised image is utilized to recover the noised patch. One drawback of their work is the extensive duration required for inference. Inspired by their approach, Iqbal et al. (2023) introduced masking-based regularization to train diffusion models, exploring various masking techniques applied at different levels of image modeling and frequency modeling. Nevertheless, their method also leads to increased inference time. Despite the potential of these methods for reconstruction fidelity, existing approaches struggle to preserve detailed characteristics of brain structure. To tackle this, we propose conditioning the denoising process of diffusion models with a latent representation encoding information about the common patterns of the input image, thereby preserving common information in the reconstructed output.
@@ 3.1 Diffusion Models
Since the reverse process $q(x_{t-1}|x_{t})$ is intractable, it is used a DNN to approximate the distribution $p_{\theta}(x_{t-1}|x_{t})$, where $\theta$ represents the weights and biases of the network. The reverse process is then modeled using a Gaussian distribution of the form:
@@
$x_{t-1}=\sqrt{\alpha_{t}-1}\left(\frac{x_{t}-\sqrt{1-\alpha_{t}}\epsilon_{\theta}(x_{t},t)}{\sqrt{\alpha_{t}}}\right)+\sqrt{1-\alpha_{t-1}}\epsilon_{\theta}(x_{t},t)+\sigma_{t}\epsilon_{t},$ (4)
where $\epsilon_{t}\sim\mathcal{N}(0,I)$ is standard Gaussian noise. For training the decoder, Ho et al. (2020) reformulated the variational lower bound objective function and considered the objective of predicting the total noise component added to the original image to create the noisy image at a given step. The loss function is then given by the squared difference between the predicted noise $\epsilon_{\theta}(x_{t},t)$ and the actual noise $\epsilon_{t}$, for a given time step $t$, using a U-Net Ronneberger et al. (2015):
@@
By setting $\sigma_{t}=0$ in Equation (4), the coefficient before the random noise $\epsilon_{t}$ becomes zero, resulting in the following deterministic process of generating a new sample $x_{t}$. This modification allows to speed up the sampling process without degrading the quality of the generated samples Bishop and Bishop (2024).
@@ 3.2 Contrastive Learning
Contrastive Learning (CL) approaches aim at pulling positive samples’ representations (e.g. of the same class) closer together while repelling representations of negative ones (e.g. different classes) apart from each other. Contrasting positive pairs against negative ones is an idea that dates back to previous research (Hadsell et al., 2006; Oord et al., 2019; Tian et al., 2020) and has seen various applications in different tasks, such as face recognition (Schroff et al., 2015). Let $x\in\mathcal{X}$ be an anchor sample, $x^{+}$ a positive sample (wrt to the anchor), and $x^{-}$ a negative sample. CL methods look for a parametric mapping function $f:\mathcal{X}\rightarrow\mathbb{S}^{d-1}$ that maps semantically similar samples close together in the representation space (i.e. a hypersphere) and dissimilar samples far away from each other. Once pre-trained, $f$ is fixed, and its representation is evaluated on a downstream task, such as classification, through linear evaluation on a test set. Depending on how positive and negative samples are defined, CL can be employed in self-supervised Chen et al. (2020) or supervised Khosla et al. (2020) settings.
@@ 4 Method
4.1 Common Features Encoder
The first block of our proposed approach is represented by the common features encoder. This encoder has the goal of learning an input representation which is invariant to the target variable (e.g. presence of eyeglasses in facial images or tumor on brain MRIs) but retains the common information of the input sample. The rationale of this approach is that an invariant representation can allow us to correctly reconstruct a realistic normal version of the input image, as it only encodes its common information. To explain the formulation of our encoder, we introduce two definitions specifying the properties that ensure that the encoder preserves the common information of the image.
@@ 4.1 Common Features Encoder
where $\epsilon\geq 0$. As the margin $\epsilon$ increases, $f$ will provide a better separation between different subjects. In practice, Eq. 6 can be expressed in terms of cosine similarity [^1] As representations are normalized, i.e. $||f(x)||_{2}=1$, then $sim(f(x),f(x^{+}))=1-d(f(x),f(x^{+}))$ where $d$ is a L2-distance function.: $sim(f(x),f(x^{-}_{j}))-sim(f(x),f(x^{+}))\leq-\epsilon\quad\forall j$ which corresponds to the $\epsilon$-InfoNCE loss Barbano et al. (2023):
@@
where $s^{+}$ and $s^{-}_{j}$ are shorthand notations for $sim(f(x),f(x^{+}))$ and $sim(f(x),f(x^{-}_{j})$ respectively. To obtain a sample $x^{+}$ of the same subject of $x$, if it is not available in the training data, it is possible to employ an augmentation scheme such as in SimCLR Chen et al. (2020), i.e. $x^{+}=t(x)$ where $t\sim\mathcal{T}$ is an augmentation operator sampled from a family of standard augmentation $\mathcal{T}$ (e.g. random transformations, cropping, etc.).
@@
(Class-invariant encoder) Denoting with $\mathcal{H}\subset\mathcal{X}$ the set of samples which share the same target attribute value (e.g. healthy), and assuming a binary case for simplicity, we say that $f$ is class-invariant, if $x\in\mathcal{H}$ and $x^{+}\in\mathcal{X}\setminus\mathcal{H}$. This means that the alignment in the latent space will be performed between samples with a different target attribute, hence achieving invariance. In a SSL setting, to avoid the dependence on anomalous samples, we leverage data augmentation and data manipulation techniques Dufumier et al. (2023) for approximating the distribution of the target attribute. For example, the appearance of tumors in brain MRIs can be approximated by employing random cutout, or Gaussian noise Behrendt et al. (2024) (details in Appendix B.2).
@@
(Common Features Encoder) An encoder $f$ preserves common patterns of the image if it is both instance-aligned and class-invariant.
The above definitions provide the theoretical support of the contrastive learning approach used for training the common features encoder. Considering that the learning process of the encoder is based on a contrastive learning strategy that is instance-aligned and class-invariant, our encoder is expected to produce features which preserve common information of the input image, regardless of whether the image belongs to the background or target set.
@@ 4.2 Conditional Diffusion-based Decoder
Our conditional diffusion decoder $p_{\theta}(x_{t-1}|x_{t},z_{bg})$ takes as input the noisy sample $x_{t}$ and the common feature $z_{bg}\in\mathbb{R}^{1\times d}=f_{\theta}(x_{0})$, a non-spatial vector of dimension $d$ that encodes common patterns observed in the input sample, derived from the properties of the common features encoder $f_{\theta}$ (Section 4.1). Our primary objective is to reconstruct the background version of the input image, preserving its common information. Hence, we deviate from the step-wise sampling process in Equation (3), which is typically employed to generate new images from noise $p_{\theta}(x_{T:1}),x_{T}\sim\mathcal{N}(0,I)$. Instead, we directly estimate the input image $\hat{x_{0}}\sim p_{\theta}(x_{0}|x_{t},z_{bg})$ given the noisy sample $x_{t}$ (more details in Algorithm 2 in Appendix A.2). This is achieved by revising Equation (2), following Song et al. (2021), enabling to prediction of the denoised observation, which is an estimation of $x_{0}$ given $x_{t}$:
@@
Then, the model is trained (Algorithm 1 in Appendix A.1) by minimizing the following objective function Preechakul et al. (2022), which is a modified version of the MSE objective in Equation (5):
@@
We utilize adaptive group normalization layers (AdaGN) to condition the U-Net, as introduced in prior works Dhariwal and Nichol (2021); Preechakul et al. (2022). AdaGN incorporates the timestep and conditional variable embedding into each residual block by applying channel-wise scaling and shifting on the normalized intermediate activations $h\in\mathbb{R}^{c\times h\times w}$: $\textrm{AdaGN}(h,t,z_{bg})=z_{i}(t_{s}\textrm{GroupNorm}(h)+t_{b})$, where $z_{i}\in\mathbb{R}^{c}=\textrm{Affine}(z_{bg})$ and $(t_{s},t_{b})\in\mathbb{R}^{2\times c}=MLP(\psi(t))$ is the output of a multilayer perceptron incorporating a sinusoidal encoding function $\psi$. Additional details are provided in Appendix A.
5 Experiments
5.1 Experimental Data

### v3

# Unsupervised contrastive analysis for anomaly detection in brain MRIs via conditional diffusion models

@@ Abstract
Contrastive Analysis (CA) detects anomalies by contrasting patterns unique to a target group (e.g., unhealthy subjects) from those in a background group (e.g., healthy subjects). In the context of brain MRIs, existing CA approaches rely on supervised contrastive learning or variational autoencoders (VAEs) using both healthy and unhealthy data, but such reliance on target samples is challenging in clinical settings. Unsupervised Anomaly Detection (UAD) offers an alternative by learning a reference representation of healthy anatomy without the need for target samples. Deviations from this reference distribution can indicate potential anomalies. In this context, diffusion models have been increasingly adopted in UAD due to their superior performance in image generation compared to VAEs. Nonetheless, precisely reconstructing the anatomy of the brain remains a challenge. In this work, we propose an unsupervised framework to improve the reconstruction quality by training a self-supervised contrastive encoder on healthy images to extract meaningful anatomical features. These features are used to condition a diffusion model to reconstruct the healthy appearance of a given image, enabling interpretable anomaly localization via pixel-wise comparison. We validate our approach through a proof-of-concept on a facial image dataset and further demonstrate its effectiveness on four brain MRI datasets, achieving state-of-the-art anomaly localization performance on the NOVA benchmark.
keywords:
Contrastive Analysis , Unsupervised Anomaly Detection , Brain MRI , Contrastive Learning , Neuroimaging , Diffusion Models
journal: Pattern Recognition Letters
[…]
[ubi] organization=Universidade da Beira Interior and NOVA LINCS,city=Covilhã, country=Portugal
[…]
[feup] organization=Faculdade de Engenharia da Universidade do Porto,city=Porto, country=Portugal
[…]
[unito] organization=University of Turin,city=Turin, country=Italy
[…]
[inesctec] organization=INESC TEC,city=Porto, country=Portugal
@@ 1 Introduction
Figure 1: Overview of the proposed contrastive-guided conditional diffusion model. The target-invariant encoder $f_{\theta}(.)$ receives the input image $x_{0}$ to encode the latent representation $z_{bg}$, which captures common information. The decoder $p_{\theta}(x_{0}|x_{t},z_{bg})$ receives the noisy sample $x_{t}$ jointly with the common feature $z_{bg}$ to reconstruct the normal version of the input image. The saliency map is obtained by $\Delta_{SM}=\left|x_{0}-\hat{x_{0}}\right|$, where high intensity values indicate the anomaly patterns. Intuitively, the class of the image can be determined based on the magnitude of $\bar{\Delta_{SM}}$.
@@
Despite substantial progress in image classification models, they still face two significant challenges: i) dependence on extensive amounts of labeled data; and ii) limited interpretability. These challenges are particularly critical in sectors such as healthcare, where explaining the decision process is essential for clinicians to trust the model outcome, and unhealthy samples are scarce and hard to obtain (Patrício et al., 2023). Contrastive Analysis (CA) represents a promising solution to this problem (Zou et al., 2013; Abid et al., 2018; Weinberger et al., 2022) by learning the fundamental generative factors that distinguish a target (TG) dataset from a background (BG) dataset (referred to as anomalous patterns) and are shared across both datasets (referred to as common patterns). However, state-of-the-art CA methodologies have several limitations: i) they often result in blurred images when utilizing Variational Autoencoders (VAEs) or are susceptible to mode collapse and unstable training when using Generative Adversarial Networks (GANs) (Schlegl et al., 2019; Carton et al., 2024); ii) they depend on the availability of healthy (BG) and unhealthy (TG) samples during training; iii) they struggle to preserve common patterns of the original image during the reconstruction process. On the other hand, UAD has emerged as an established approach for modeling the distribution of healthy anatomy and identifying deviations from this distribution as potential anomalies, thus removing the dependence on annotated unhealthy (TG) samples. While Denoising Diffusion Probabilistic Models (DDPM) (Ho et al., 2020) are commonly adopted as generative models in UAD, their training process lacks explicit constraints to preserve common patterns during reconstruction, hindering the fidelity of normal (healthy) reconstructions. Throughout this paper, we use the terms CA and UAD interchangeably, when referring to the task of detecting anomalous patterns in images.
In this work, we propose an unsupervised methodology for the task of anomaly detection in brain MRIs. Specifically, we build upon a contrastive-guided conditional diffusion model (Figure 1) to reconstruct a healthy version of the input image while preserving its common anatomical patterns. This is achieved by conditioning the denoising diffusion process on a target-invariant latent feature. To this end, we employ a self-supervised contrastive encoder (referred to as target-invariant encoder) trained exclusively on samples from the healthy (BG) dataset. To approximate potential anomalous variation factors, we apply data augmentation techniques such as random cutout and Gaussian noise, which simulate features like tumor presence in brain MRIs. During inference, the diffusion model processes unseen samples belonging to the healthy (BG) or unhealthy (TG) distributions by reconstructing the input image such that the anomalous patterns are replaced with learned common (healthy) patterns. By computing pixel-wise differences between the original image and its generated counterpart, high-intensity values can be used to i) classify the input image as background (normal) or target (anomalous), based on the magnitude of the reconstruction error; and ii) localize potential anomalies, as illustrated in Figure 1 (bottom-right).
Proof-of-concept validation on facial imagery data demonstrates the utility of our method. Additionally, an extensive comparative analysis against state-of-the-art methods for UAD on four brain MRI benchmarks highlights the effectiveness of our approach, achieving top performance on the recent and challenging evaluation-only NOVA benchmark (Bercea et al., 2025). Our major contributions are as follows:
  • We propose a methodology for learning common features between background (healthy) and target (unhealthy) distributions, enabling the generation of a healthy (normal) counterpart of the input image by encoding only common information;
  • We introduce a target-invariant encoder that learns representations which are both class-invariant and instance-aligned, facilitating common features preservation during the generative process;
  • We conduct a comprehensive evaluation of our approach on a facial imagery dataset and four brain MRI datasets, encompassing both healthy (background) and tumor (target) images, to validate the generalization of our approach;
  • Our method achieves state-of-the-art performance in anomaly localization on the challenging NOVA benchmark.
@@ 2 Related Work
Early work in contrastive analysis (CA) relied on the use of Variational Autoencoders (VAEs) (Baur et al., 2021; Behrendt et al., 2022; Louiset et al., 2023). A recent method proposed by Louiset et al. (2023), SepVAE, aims to distinguish common (healthy) from class-specific (unhealthy) patterns in image data. They utilized VAEs with regularization to encourage disentanglement between common and salient representations, along with a classification loss to separate target and background salient factors. However, the resulting reconstructions were often blurry, limiting their interpretability and utility. Alternative approaches (Schlegl et al., 2019; Carton et al., 2024) leveraging Generative Adversarial Networks (GANs) have been explored, but they suffer from issues like mode collapse and unstable training. More recently, Diffusion Denoising Probabilistic Models (Song et al., 2021; Preechakul et al., 2022) (DDPMs) have emerged as a promising alternative for high-quality image generation, addressing the drawbacks of both GANs and VAEs. Our work aligns closely with the setting of CA and UAD. In the context of brain MRI, several works have been proposed (Behrendt et al., 2024; Iqbal et al., 2023; Behrendt et al., 2025). For example,  Behrendt et al. (2024) introduced a patch-based diffusion model for UAD in brain MRIs. They divide the input image into predefined patches and apply noise to each patch individually in the forward pass. In the backward pass, the partly noised image is utilized to recover the noised patch. One drawback of their work is the extensive duration required for inference. More recently, Behrendt et al. (2025) proposed to conditioning a DDPM to better capture intensity characteristics and domain shifts in brain MRIs during reconstruction, thereby improving segmentation performance. In contrast, our method incorporates contrastive learning during encoder training to explicitly enforce a target-invariant embedding space.
@@ 3.1 Diffusion Models
Since the reverse process $q(x_{t-1}|x_{t})$ is intractable, a DNN is used to approximate the distribution $p_{\theta}(x_{t-1}|x_{t})$, where $\theta$ represents the weights and biases of the network. The reverse process is then modeled using a Gaussian distribution of the form:
@@
[…]
where $\epsilon_{t}\sim\mathcal{N}(0,I)$ is standard Gaussian noise. For training the decoder, Ho et al. (2020) reformulated the variational lower bound objective function and considered the objective of predicting the total noise component added to the original image to create the noisy image at a given step. The loss function is then given by the squared difference between the predicted noise $\epsilon_{\theta}(x_{t},t)$ and the actual noise $\epsilon_{t}$, for a given time step $t$, using a U-Net (Ronneberger et al., 2015):
@@
By setting $\sigma_{t}=0$ in Equation (4), the coefficient before the random noise $\epsilon_{t}$ becomes zero, resulting in the following deterministic process of generating a new sample $x_{t}$. This modification allows to speed up the sampling process without degrading the quality of the generated samples (Bishop & Bishop, 2024).
@@ 3.2 Contrastive Learning
Contrastive Learning (CL) approaches aim at pulling positive samples’ representations (e.g. of the same class) closer together while repelling representations of negative ones (e.g. different classes) apart from each other. Contrasting positive pairs against negative ones is an idea that dates back to previous research (Hadsell et al., 2006; Oord et al., 2019; Tian et al., 2020) and has seen various applications in different tasks, such as face recognition (Schroff et al., 2015). Let $x\in\mathcal{X}$ be an anchor sample, $x^{+}$ a positive sample (wrt the anchor), and $x^{-}$ a negative sample. CL methods look for a parametric mapping function $f:\mathcal{X}\rightarrow\mathbb{S}^{d-1}$ that maps semantically similar samples close together in the representation space (i.e. a hypersphere) and dissimilar samples far away from each other. Once pre-trained, $f$ is fixed, and its representation is evaluated on a downstream task, such as classification, through linear evaluation on a test set. Depending on how positive and negative samples are defined, CL can be employed in self-supervised (Chen et al., 2020) or supervised (Khosla et al., 2020) settings.
@@ 4 Method
Our method employs a target-invariant encoder to extract a latent representation capturing common features of the input image. A diffusion-based model is then conditioned on this latent representation to reconstruct a normal (healthy) counterpart of the input. The absolute difference between the reconstructed and original images reveals potential anomalies, where brighter areas indicate abnormalities. An overview of the proposed approach is presented in Figure 1.
4.1 Target-Invariant Encoder
The first block of our proposed approach is represented by the target-invariant encoder. This encoder has the goal of learning an input representation which is invariant to the target variable (e.g. presence of eyeglasses in facial images or tumor in brain MRIs) but retains the common information of the input sample. Figure 1 (top-right) presents a visualization of the embedding space after projecting both healthy and unhealthy samples from the trained encoder. The rationale of this approach is that an invariant representation can allow us to correctly reconstruct a realistic normal version of the input image, as it only encodes its common (healthy) information. To explain the formulation of our encoder, we introduce three definitions specifying the properties that ensure that the encoder preserves the common information of the image.
@@ 4.1 Target-Invariant Encoder
where $\epsilon\geq 0$. As the margin $\epsilon$ increases, $f$ will provide a better separation between different subjects. In practice, Eq. 6 can be expressed in terms of cosine similarity [^1] As representations are normalized, i.e. $||f(x)||_{2}=1$, then $sim(f(x),f(x^{+}))=1-d(f(x),f(x^{+}))$ where $d$ is a L2-distance function.: $sim(f(x),f(x^{-}_{j}))-sim(f(x),f(x^{+}))\leq-\epsilon\quad\forall j$ which corresponds to the $\epsilon$-InfoNCE loss (Barbano et al., 2023):
@@
where $s^{+}$ and $s^{-}_{j}$ are shorthand notations for $sim(f(x),f(x^{+}))$ and $sim(f(x),f(x^{-}_{j}))$ respectively. To obtain a sample $x^{+}$ of the same subject of $x$, if it is not available in the training data, it is possible to employ an augmentation scheme such as in SimCLR (Chen et al., 2020), i.e. $x^{+}=t(x)$ where $t\sim\mathcal{T}$ is an augmentation operator sampled from a family of standard augmentation $\mathcal{T}$ (e.g. random transformations, cropping, etc.).
@@
(Class-invariant encoder) Denoting with $\mathcal{H}\subset\mathcal{X}$ the set of samples which share the same target attribute value (e.g. healthy), and assuming a binary case for simplicity, we say that $f$ is class-invariant, if $x\in\mathcal{H}$ and $x^{+}\in\mathcal{X}\setminus\mathcal{H}$. This means that the alignment in the latent space will be performed between samples with a different target attribute, hence achieving invariance. In a SSL setting, to avoid the dependence on anomalous samples, we leverage data augmentation and data manipulation techniques (Dufumier et al., 2023) for approximating the distribution of the target attribute. For example, the appearance of tumors in brain MRIs can be approximated by employing random cutout, or Gaussian noise (Behrendt et al., 2024) (more details in the supplementary material).
@@
(Target-Invariant Encoder) An encoder $f$ preserves common patterns of the image if it is both instance-aligned and class-invariant.
The above definitions provide the theoretical support of the contrastive learning approach used for training the target-invariant encoder. Considering that the learning process of the encoder is based on a contrastive learning strategy that is instance-aligned and class-invariant, our encoder is expected to produce features which preserve common information of the input image, regardless of whether the image belongs to the background or target set.
@@ 4.2 Conditional Diffusion-based Decoder
Our conditional diffusion decoder $p_{\theta}(x_{t-1}|x_{t},z_{bg})$ takes as input the noisy sample $x_{t}$ and the common feature $z_{bg}\in\mathbb{R}^{1\times d}=f_{\theta}(x_{0})$, a non-spatial vector of dimension $d$ that encodes common patterns observed in the input sample, derived from the properties of the target-invariant encoder $f_{\theta}$ (Section 4.1). Our primary objective is to reconstruct the background version of the input image, preserving its common information. Hence, we deviate from the step-wise sampling process in Equation (3), which is typically employed to generate new images from noise $p_{\theta}(x_{T:1}),x_{T}\sim\mathcal{N}(0,I)$. Instead, we directly estimate the input image $\hat{x_{0}}\sim p_{\theta}(x_{0}|x_{t},z_{bg})$ given the noisy sample $x_{t}$ (more details in Algorithm 2 in the supplementary material). This is achieved by revising Equation (2), following (Song et al., 2021), allowing prediction of the denoised observation, which is an estimation of $x_{0}$ given $x_{t}$:
@@
The model is then trained (Algorithm 1 in the supplementary material) by minimizing the following objective function (Preechakul et al., 2022), which is a modified version of the MSE objective in Equation (5):
@@
We condition the U-Net using adaptive group normalization (AdaGN), which integrates timestep and conditioning embeddings into each residual block via channel-wise scaling and shifting of normalized activations. Further details are provided in the supplementary material.
5 Experimental Setup
5.1 Datasets